# Your Student is Better Than Expected: Adaptive Teacher-Student Collaboration for Text-Conditional Diffusion Models

### v2

@@ 3.1 Delving deeper into the student performance
Student-teacher similarity. First, we evaluate the student’s ability to imitate the teacher. We compute pairwise similarities between the student and teacher images generated for the same text prompts and initial noise. As a similarity measure, we consider DreamSim [9] tuned to align with the human perception judgments. For evaluation, we consider 5000 prompts from the COCO2014 [22] validation split.
Primarily, we observe that many student samples are highly distinct from the teacher ones. A few image pairs are presented in Figure 3a. Figure 3c presents the human vote distribution for three distance ranges: low, medium, and high. Interestingly, most of the student wins appear when its samples are highly different from the teacher ones. This brings us to our second observation:
@@
Also, we evaluate the relative gap in sample quality against the similarity between the teacher and student outputs. To measure the quality of individual samples, we use ImageReward [53], which shows a positive correlation with human preferences in terms of image fidelity and prompt alignment. The gap in sample quality is calculated as the difference between the ImageReward scores for student and teacher samples. We observe that highly distinct samples likely have a significant difference in quality. Importantly, this holds for both student failures and successes, as shown in Figure 3b. Therefore, effectively detecting the positive student samples and improving the negative ones can potentially increase the generative performance.
@@ 3.2 Method
Then, comparing the scores of the teacher and student samples, one can conclude which one is better. However, in practice, we avoid expensive teacher inference to preserve the efficiency of our approach. Therefore, a decision must be made having access only to the student sample $\mathcal{X}^{\mathbf{S}}$. To address this problem, we introduce a cut-off threshold $\tau$ which is a $k$-th percentile of the IR score tuned on a hold-out subset of student samples. The details on the threshold tuning are described in C. During inference, the IR score is calculated only for $\mathcal{X}^{\mathbf{S}}$. If it exceeds the threshold $\tau$, we accept the sample and avoid further teacher involvement. Interestingly, we observe that it is often possible to reproduce the accuracy of the full-reference estimation by varying $\tau$ (see Figure 7). Also, note that IR calculation costs are negligible compared to a single diffusion step, see D.6.
@@ 4.1 Text-guided image synthesis
Also, we evaluate user preferences using side-by-side comparisons conducted by professional assessors. We select $600$ random text prompts from the COCO2014 validation set and $600$ from LAION-Aesthetics [^2] https://laion.ai/blog/laion-aesthetics/. More details on the human evaluation pipeline are provided in D.1.
@@
Figure 12: User preferences for different accuracy levels of the no-reference decision-making procedure. Current state represents the results using ImageReward. The results for higher accuracy rates demonstrate the future gains if the oracle performance improves.
@@
Effect of oracle accuracy. The potential bottleneck of our approach is a poor correlation of existing text-to-image automated estimators with human preferences. For example, ImageReward usually exhibits up to $65\%$ agreement with annotators. Moreover, it remains unclear what oracle accuracy can be achieved with no-reference decision-making, even if the estimator provides the perfect agreement.
In Figure 12, we conduct a synthetic experiment examining the effect of the oracle accuracy on our scheme performance to reveal its future potential. We compare the adaptive refinement method ($10$ steps) to SD1.5 ($50$ steps) manually varying the oracle accuracy. We observe significant future gains even for the $75\%$ accuracy rate.
@@
4.2 Distribution diversity
Figure 13: Diversity human scores collected for different methods.
CD-SDXL demonstrates significantly better diversity than ADD-XL but still produces less various images compared to the SDXL teacher. CD-SD1.5 performs similarly to the SD1.5 teacher. Also, both adaptive strategies increase the diversity of the SDXL student models, especially the regeneration one. We also provide a few samples generated with different distilled models in D.5.
4.3 Text-guided image editing
@@
Table 1: Comparison of SDEdit using different approaches in terms of reference preservation DINOv2 and editing quality (ImageReward) for the strength $0.6$.
@@ 4.4 Controllable image generation
For the teacher model, the default ControlNet sampling configuration is used: $20$ sampling steps of the UniPC [58] solver. In our adaptive approach, we use the refinement strategy with $10$ steps of the same solver. For performance evaluation, we conduct the human preference study for each task on $600$ examples and provide more details in D.4.
@@ 5 Conclusion
This work investigates the performance of the distilled text-to-image models and demonstrates that they may consistently outperform the teachers on many samples. We design an adaptive text-to-image generation pipeline that takes advantage of successful student samples and, in combination with the teacher model, outperforms other alternatives for low and high inference budgets. Importantly, text-to-image automated estimators and distillation methods still have room for further development. Therefore, we believe the proposed approach may be even in higher demand when future works make a significant step forward in these directions.
@@ B.2 Other distilled text-to-image models
$\hat{\mathcal{O}}=\mathbf{S}(\mathcal{I})$   // Student prediction
@@ D.3 Experimental setup (SDXL)
D.4 Controllable generation
For both tasks, we use the adaptive refinement strategy and set the rollback value $\sigma$ to $0.5$. We perform $5$ steps for the student generation and $10$ steps for the refinement with the UniPC solver. The cut-off thresholds correspond to $70$ and $50$ ImageReward percentiles for the mask-guided and edge-guided generation, respectively. We select random $600$ image-text pairs from the COCO2014 validation set for the edge-guided generation. For the mask-guided generation, we use $600$ semantic segmentation masks from the ADE20K dataset [59] and use the category names as the text prompts. To evaluate the performance, we conduct the human study similar to D.1.
@@ D.5 Distribution diversity
In Figure 27, we illustrate the diversity of images generated with different distilled text-to-image models (ADD-XL, CD-SDXL and CD-SD1.5). Each column corresponds to a different initial noise (seed). We notice that ADD-XL exhibits the lowest diversity compared to the CD-based counterparts.
D.6 ImageReward inference costs
@@
D.7 Additional visualizations
In Figure 23, we present more examples where the student outperforms its teacher according to the human evaluation. Figures 24, 25, 26 provide more qualitative comparisons of our approach for different tasks.

### v3

@@ 3.1 Delving deeper into the student performance
The student can surpass its teacher in a substantial portion of image samples.
@@
Student-teacher similarity. First, we evaluate the student’s ability to imitate the teacher. We compute pairwise distances between the student (S) and teacher (T) images generated for the same text prompts and initial noise. As a distance measure, we use DreamSim [9] tuned to be aligned with the human perception judgments. For evaluation, we consider 5000 prompts from the COCO2014 [22] validation split.
Primarily, we observe that many student samples are highly distinct from the teacher ones. A few image pairs are presented in Figure 3a. Figure 3c presents the human vote distribution for low ($0{-}20\%$), medium ($40{-}60\%$) and high ($80{-}100\%$) distance ranges. Interestingly, most of the student wins appear when its samples are highly different from the teacher ones. This brings us to our second observation:
@@
Also, we evaluate the relative gap in sample quality against the similarity between the teacher and student outputs. To measure the quality of individual samples, we use ImageReward [53], which shows a positive correlation with human preferences in terms of image fidelity and prompt alignment. The divergence in quality is calculated as the difference between the ImageReward scores for student and teacher samples. We observe that highly distinct samples likely have a significant difference in quality. Importantly, this holds for both student failures and successes, as shown in Figure 3b. Therefore, effectively detecting the positive student samples and improving the negative ones can potentially increase the generative performance.
@@ 3.2 Method
Then, comparing the scores of the teacher and student samples, one can conclude which one is better. However, in practice, we avoid expensive teacher inference to preserve the efficiency of our approach. Therefore, a decision must be made having access only to the student sample $\mathcal{X}^{\mathbf{S}}$. To address this problem, we introduce a cut-off threshold $\tau$ which is a $k$-th percentile of the IR score tuned on a hold-out subset of student samples. The details on the threshold tuning are described in C. During inference, the IR score is calculated only for $\mathcal{X}^{\mathbf{S}}$. If it exceeds the threshold $\tau$, we accept the sample and avoid further teacher involvement. Interestingly, we observe that it is often possible to reproduce the accuracy of the full-reference estimation by varying $\tau$ (see Figure 7). Also, note that IR calculation costs are negligible compared to a single diffusion step, see D.7.
@@ 4.1 Text-guided image synthesis
Also, we evaluate user preferences using side-by-side comparisons conducted by professional assessors. We select $600$ random text prompts from the COCO2014 validation set and $600$ from LAION-Aesthetics. More details on the human evaluation pipeline are provided in D.1.
@@
In D.5, we also investigate how our approach affects the distribution diversity. D.4 aims to reveal the potential gains of our approach if the oracle accuracy increases in the future.
@@
Table 1: Comparison of SDEdit using different approaches in terms of reference preservation and editing quality for the strength $0.6$.
4.2 Text-guided image editing
@@ 4.3 Controllable image generation
For the teacher model, the default ControlNet sampling configuration is used: $20$ sampling steps of the UniPC [58] solver. In our adaptive approach, we use the refinement strategy with $10$ steps of the same solver. For performance evaluation, we conduct the human preference study for each task on $600$ examples and provide more details in D.6.
@@ 5 Conclusion
This work investigates the performance of the distilled text-to-image models and demonstrates that they may consistently outperform the teachers on many samples. We design an adaptive text-to-image generation pipeline that takes advantage of successful student samples and, in combination with the teacher model, outperforms other alternatives for low and high inference budgets.
@@ B.2 Other distilled text-to-image models
$\hat{\mathcal{O}}=\mathbf{S}(\mathcal{I})$
  // Student prediction
@@ Appendix C Cut-off threshold tuning
We also show that hyperparameter tuning is straightforward and requires negligible effort. To verify this, we tune the threshold using the various number of prompts (Tab. 2). We can see that $500$ prompts are sufficient for the threshold convergence. Thus, the process only needs to generate $500$ student samples and takes ${\sim}2.5$ minutes.
Prompts $150$ $250$ $500$ $5000$ $25000$
[…]
Table 2: The threshold value tuned using prompts from Diffusion-DB and PickScore datasets and the time required for tuning.
@@ D.3 Experimental setup (SDXL)
D.4 Effect of oracle accuracy
Figure 16: User preferences for different accuracy levels of the no-reference decision-making procedure. the automated sample estimator. Current state represents the results using ImageReward. The results for higher accuracy rates demonstrate the future gains if the oracle performance improves.
The potential bottleneck of our approach is a poor correlation of existing text-to-image automated estimators with human preferences. For example, ImageReward usually exhibits up to $65\%$ agreement with annotators. Moreover, it remains unclear what oracle accuracy can be achieved with no-reference decision-making, even if the estimator provides the perfect agreement. In Figure 16, we conduct a synthetic experiment examining the effect of the oracle accuracy on our scheme performance to reveal its future potential. We compare the adaptive refinement method ($10$ steps) to SD1.5 ($50$ steps) manually varying the oracle accuracy. We observe significant future gains even for the $75\%$ accuracy rate.
@@ D.5 Distribution diversity
Figure 17: Diversity human scores collected for different methods.
CD-SDXL demonstrates significantly better diversity than ADD-XL but still produces less various images compared to the SDXL teacher. CD-SD1.5 performs similarly to the SD1.5 teacher. Also, both adaptive strategies increase the diversity of the SDXL student models, especially the regeneration one. In Figure 30, we illustrate the diversity of images generated with different distilled text-to-image models (ADD-XL, CD-SDXL and CD-SD1.5). Each column corresponds to a different initial noise (seed). We notice that ADD-XL exhibits the lowest diversity compared to the CD-based counterparts.
D.6 Controllable generation
For both tasks, we use the adaptive refinement strategy and set the rollback value $\sigma$ to $0.5$. We perform $5$ steps for the student generation and $10$ steps for the refinement with the UniPC solver. The cut-off thresholds correspond to $70$ and $50$ ImageReward percentiles for the mask-guided and edge-guided generation, respectively. We select random $600$ image-text pairs from the COCO2014 validation set for the edge-guided generation. For the mask-guided generation, we use $600$ semantic segmentation masks from the ADE20K dataset [59] and use the category names as the text prompts. For evaluation, we conduct the human study similar to D.1.
D.7 ImageReward inference costs
@@
D.8 Additional visualizations
In Figure 23, 24 we present qualitative verification of the first observation (the student sometimes outperforms its teacher according to the human evaluation). Figure 25 supports the second observation (the student wins are more likely where its samples differ from the teacher ones). In Figure 26 we demonstrate two adaptive strategies (refining and regeneration), the examples confirm that the refinement strategy improves the image fidelity and does not significantly alter the textual alignment, while the regeneration strategy may improve textual alignment.
Figures 27, 28, 29 provide more qualitative comparisons of our approach for different tasks.
@@ D.8 Additional visualizations
Figure 24: Additional examples where the student (CD-SDXL) outperforms its teacher (SDXL) according to the human evaluation.
Figure 25: Visualization of the student (CD-SDXL) and teacher (SDXL) samples for low and high distance ranges. The green outline corresponds to wins, while the red one - losses.
Figure 26: Visual examples of the refining and regeneration adaptive strategies.